If the half of a reactant in a 1st order reaction is 25 minutes, then what is the extent of reaction after 2 minutes?

1
Expert's answer
2021-05-31T02:08:45-0400

Ao = Initial amount of reactant.

A = Un-reacted amount of reactant.

K= Rate constant

t = time consumed

t"\frac{1}{2}" = Half-Life of the reaction that is the time taken for half of the reactant to react.

After 2mins, 0.945part of the initial reactant concentration remains unreacted and 0.055part of the reactant has reacted. So nearly, 5.5% of the reaction has occurred.

In"\frac{A^0}{A}=Kt"

For half of the reaction, K"=\frac{0.693}{t^1/2}"


"=\frac{0.693}{25}=0.028"

For 2minutes in"\frac{A^0}{A}=0.028×2=0.056"


"\frac{A^0}{A}=1.058"


A=0.945Ao